# Dynamic Correction of Erroneous State Estimates via Diffusion Bayesian Exploration

### v2

@@ 1 Introduction
In disaster emergency management, early decisions play a pivotal role in shaping the outcome of a crisis steel2010bayesian; kruschke2010bayesian. An initial state estimate—for example, an early guess of a hazardous leak’s location—is often made with scant data under intense time pressure. Such early estimates carry great weight in guiding subsequent actions, yet they are highly vulnerable to error due to uncertainty and human subjective judgment. If the initial assumption is mistaken, the entire response effort can become locked into a false premise, failing to adjust even as new data arrives. This lock-in effect in high-stakes scenarios leads to catastrophic delays, misallocation of resources, and heightened risks to society. The fragility of such initial state assumptions thus poses a serious challenge for social decision-making processes, revealing the need for methods that can rapidly correct initial mistakes in real time.
@@
From a robotics and Bayesian inference perspective, the above scenario is essentially a state estimation problem: an agent must form an initial belief (prior) about the latent state (e.g., the hazard’s location) and update this belief as new sensor data arrives. If the initial belief is mis-specified, the goal is to correct this incorrect state estimate in real time by incorporating incoming evidence. For clarity, we use the term initial policy error to denote such a mistaken initial state estimate under prior uncertainty, which our approach aims to correct. In recent years, particle filtering (PF) has become a go-to approach for sequential Bayesian state estimation in these settings ristic2013particle; gordon1993novel; doucet2001introduction. PF offers a principled way to update beliefs by fusing sensor data with prior knowledge, making it a representative framework to tackle the problem of erroneous initial state estimates in emergency response and other high-stakes decision domains. However, existing Bayesian filtering methods fox2003bayesian; smidl2008variational struggle with erroneous initial state estimates, often failing to recover when the true state lies outside the initial belief. 1) Bootstrap particle filters candy2007bootstrap; gordon1995bayesian tend to remain confined to the support of the initial prior, hindering exploration beyond the originally assumed region when operated under a stationary bootstrap baseline with zero transition and no rejuvenation. We formalize this baseline-specific lock-in as Stationarity-Induced Posterior Support Invariance (S-PSI): under the zero-transition, no-rejuvenation baseline, the posterior cannot escape the initial support. Importantly, S-PSI is not an inherent limitation of PF; classical countermeasures (e.g., jittering, roughening) can in principle expand support (we include these as baselines in our experiments). 2) More advanced PF variants,such as auxiliary particle filters mountney2009bayesian; branchini2021optimized or filters with optimal proposal distributions, can partially mitigate bias when the true state has a small non-zero prior probability fox2001kld; douc2005comparison; liu1998sequential; doucet2000sequential; arulampalam2002tutorial, but they fundamentally cannot handle cases where the true state was assigned zero initial probability. Under the standard Bayesian update, any region with zero prior mass will remain at zero posterior mass indefinitely, meaning the algorithm can never discover a completely excluded possibility under this baseline. 3) Attempts to address these issues by augmenting particle filters have had limited success. Some works inject noise or broaden the prior artificially, and others integrate reinforcement learning (RL) with PF shi2024autonomous; zhao2022deep; park2022source to actively guide sensor exploration. While such approaches can improve data collection, they may inherit the same blind spots from a mis-specified prior and often introduce significant complexity and resource demands. Without a new perspective, Bayesian trackers and decision methods remain at risk of locking onto an incorrect initial belief, especially under the S-PSI.
@@ 2 Related Work
In emergency localization scenarios wu2021research; hite2019bayesian, Bayesian filtering fox2003bayesian; smidl2008variational; Quinlan2010Multiple methods such as the bootstrap particle filter leverage Bayesian inference to iteratively update state estimates but typically assume correctly-specified initial priors, thus may become ineffective under severely misaligned early assumptions when operated under a stationary bootstrap baseline (zero transition, no rejuvenation), due to Stationarity-Induced Posterior Support Invariance (S-PSI) gordon1995bayesian; candy2007bootstrap. S-PSI is a baseline-specific diagnostic rather than an inherent limitation of PF; classical countermeasures such as jittering/roughening or resample–move can, in principle, expand support, and we include them as baselines in §5. Advanced particle filter variants, including auxiliary particle filters mountney2009bayesian; branchini2021optimized and filters using optimal proposal distributions, improve proposal quality and sample efficiency but still fail when the initial belief assigns zero prior probability to the true state fox2001kld; arulampalam2002tutorial, i.e., in the presence of a zero-prior barrier without explicit support expansion. Meanwhile, information-theoretic methods like Infotaxis vergassola2007infotaxis, Entrotaxis hutchinson2018entrotaxis and DCEE chen2021dual focus sensor motions on maximizing expected information gain or reducing entropy, yet they typically operate within the belief support induced by the initial prior and thus cannot systematically correct severe prior misalignment in real time without a support-expanding inference layer. More recently, integrated reinforcement learning and particle filtering (RL-PF) methods have emerged—e.g., AGDC and its variants using KL-divergence or entropy-based intrinsic rewards shi2024autonomous, PC-DQN zhao2022deep, and GMM-PFRL park2022source. While these RL-driven methods exhibit stronger adaptive exploration, they can still inherit the zero-prior barrier when the underlying filtering layer does not expand support (cf. the S-PSI baseline). In contrast, our proposed DEPF explicitly addresses this gap by introducing a belief-triggered, validated support-expansion mechanism that operates at the inference layer, making it complementary to proposal-improvement filters, classical perturbations (jittering/roughening/rejuvenation), and planning/RL controllers, and yielding superior robustness under severely misaligned early assumptions.
@@ 3.5 POMDP formulation for belief‑aware sensor planning
The interaction between the mobile robot and the unknown gas plume can be framed as a partially observableMarkov decision process (POMDP) $\mathcal{M}=(\mathcal{S},\mathcal{A},\Omega,T,O,R,\gamma)$, where $\mathcal{S}$ is the latent state space containing the stationary source vector $\Theta=(x_{s},y_{s},q_{s},u_{s},\phi_{s},d_{s},\tau_{s})^{\top}$; $\mathcal{A}$ is the set of motion commands that move the robot in the plane; $\Omega$ is the observation space, where an observation at time $k$ is $o_{k}=(p_{k},z_{k})$ with $p_{k}\in\mathbb{R}^{2}$ the robot position and $z_{k}\in\mathbb{R}$ the sensor voltage. The transition kernel factorises as $T((p^{\prime},\Theta^{\prime})\mid(p,\Theta),a)=T_{p}(p^{\prime}\mid p,a)\,\delta(\Theta^{\prime}-\Theta)$: robot kinematics are deterministic while the source parameters remain unchanged. The observation model is $O(o_{k}\mid(p_{k},\Theta),a_{k-1})=p(z_{k}\mid p_{k},\Theta)$, where the likelihood $p(z_{k}\mid p_{k},\Theta)$ is the mixture‑Gaussian plume sensor model of §3.2. At decision time the agent cannot access $\Theta$ directly, so it reasons with the belief $b_{k}=p(\Theta\mid z_{1:k})$ supplied by the particle filter. We therefore feed the RL policy with the augmented information state $s^{\text{RL}}_{k}=(p_{k},z_{k},b_{k})$. The reward is chosen as the expected one‑step information gain $R_{k}=\mathbb{E}_{o_{k+1}}[D_{\mathrm{KL}}(b_{k+1}\,\|\,b_{k})]$, encouraging actions that shrink posterior uncertainty, and future rewards are discounted by $\gamma\in(0,1)$. The objective is to learn a policy $\pi^{\ast}$ that maximises the expected discounted return $J=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}R_{k+t}]$, thereby steering the robot along paths that are most informative about the hidden source. The details are provided in Appendix § O.
@@ 4 Diffusion-Driven Support Range Expansion
Diffusion-Driven Validation via MCMC: To ensure consistency with the target posterior distribution, a Metropolis-Hastings acceptance criterion hastings1970monte validates the diffused particles. For each perturbed particle $\Theta_{k}^{(i)}$, the acceptance probability is: $\alpha_{i}=\frac{w_{\text{new}}^{(i)}}{w_{\text{old}}^{(i)}}\cdot\exp\left(-\frac{1}{2}\Delta\Theta_{k}^{(i)^{T}}\Sigma^{-1}\Delta\Theta_{k}^{(i)}\right).$ A uniformly sampled random variable $u_{i}\sim\mathcal{U}(0,1)$ determines whether the particle is accepted: $\Theta_{k}^{(i)}=\Theta_{k}^{(i)}-\Delta\Theta_{k}^{(i)}\cdot\mathbb{I}(\alpha_{i}<u_{i})$, where $\mathbb{I}(\alpha_{i}<u_{i})$ is the indicator function, which equals 1 when $\alpha_{i}<u_{i}$ is true, and 0 otherwise. This step ensures that the diffusion-driven expansion aligns with the posterior distribution, preserving the accuracy of the particle filter.
@@ 5.1 Evaluation Metrics and Baseline Algorithms
To evaluate our proposed approach and compare it against baseline algorithms, we use four distinct metrics: Operational Completion Efficacy (OCE), which measures how frequently emergency response missions meet their goals, with higher scores indicating better deployment effectiveness; Average Deployment Efficiency (ADE), representing the average distance traveled by response units, where shorter distances imply more efficient routing; Response Execution Velocity (REV), quantifying the time duration from deployment to task completion, with faster times signifying more efficient operations; and Localization Precision Score (LPS), assessing the accuracy of source localization by computing the average discrepancy between estimated and actual source locations, with smaller values denoting higher accuracy. Our proposed method is evaluated alongside various baseline algorithms, grouped according to their methodological foundations. The first group merges reinforcement learning with Bayesian inference and includes a single representative, AGDC shi2024autonomous, which leverages the particle-filter RL posterior and uses intrinsic rewards derived from belief updates to guide exploration. Additionally, we include two other reinforcement learning approaches, PC-DQN zhao2022deep and GMM-PFRL park2022source, which independently leverage particle filtering parameters as states for RL training. The second group integrates planning and Bayesian inference approaches, represented by Infotaxis vergassola2007infotaxis, Entrotaxis hutchinson2018entrotaxis, and DCEE chen2021dual. Finally, we include a third group of classical SMC perturbation baselines that are not subject to S-PSI constraints: PF+Jittering liu1998sequential; doucet2000sequential, PF+Roughening gordon1993novel; gordon1995bayesian, and PF+Rejuvenation hastings1970monte; doucet2000sequential. Implementation details and hyperparameter grids are given in §H and I.
@@ Appendix A Particle Filtering: Recursive Approximation of Posterior Distributions
Step 1. Sampling (Prediction): gordon1993novel Particles are sampled from an importance distribution $q(\Theta_{k}\mid\Theta_{k-1},z_{k})$, typically approximated by the state transition model $P(\Theta_{k}\mid\Theta_{k-1})$. The propagation of particles is expressed as:
@@
Step 2. Weight Update: doucet2001introduction Particle weights are updated to reflect their likelihood given the new observation. The unnormalized weights are calculated as:
@@
Step 3. Normalization of Weights: liu1998sequential Weights are normalized to maintain their probabilistic interpretation:
@@
Step 4. Resampling: doucet2000sequential To mitigate weight degeneracy (domination by a few particles), resampling occurs when the effective particle number $N_{\text{eff}}$ falls below a threshold:
@@ C.3 Entropy-Based Weight Smoothing
which can be viewed as Bayesian smoothing (doucet2001sequential; liu1998sequential). In practice, we also adopt a standard tempering scheme
@@ C.4 Kernel-Induced Stochastic Diffusion
where $LL^{\top}=\Sigma$ is the Cholesky factorization of the weighted covariance. This is equivalent to a kernel density approximation (silverman2018density), and ensures that as $N\to\infty$ the perturbed empirical measure converges to the true posterior.
@@ C.5 Metropolis–Hastings Validation
This guarantees detailed balance and convergence to the true posterior distribution (hastings1970monte).
@@ I.3 Pseudocode
Appendix J Extended Related Work
Prior Region Misalignment in Bayesian Inference: The effectiveness of particle filtering relies heavily on the alignment between the prior distribution and the true target state. Historically, misaligned priors have led to biased state estimations, a challenge highlighted by foundational works steel2010bayesian; kruschke2010bayesian. Traditional techniques, such as importance sampling ristic2013particle, operate effectively within prior boundaries but fail for states beyond the initial support. Methods like adaptive particle filters fox2001kld and hybrid resampling strategies douc2005comparison have sought to improve sample efficiency but remain limited in their ability to dynamically adapt to global misalignment.
Support Invariance under Baseline Assumptions: In this work we formalize the Stationarity-Induced Posterior Support Invariance (S-PSI) baseline pathology, extending insights from PF’s recursive limitations doucet2000sequential. Under the stationary bootstrap setting with zero transition and no rejuvenation, particles remain confined to the initial prior support gordon1993novel. This diagnostic effect has been observed in dynamic contexts such as robot localisation thrun2002probabilistic and sensor fusion candy2016bayesian, where prior misalignment impacts performance. While classical perturbation techniques (jittering, roughening, resample–move) can relax S-PSI in principle, they operate in an always-on and inefficient manner, lacking mechanisms to decide when and how to expand.
Bayesian Inference in Out-of-Distribution Problems: Out-of-distribution (OOD) problems represent a critical challenge for Bayesian inference, as they require models to generalise beyond prior knowledge. Advances such as Posterior Networks charpentier2020posterior and Bayesian OOD detection with uncertainty exposure wangbayesian have addressed OOD detection in classification tasks but are less effective in sequential estimation. Traditional PF variants such as Rao-Blackwellised filters li2004estimation and model-switching approaches moradkhani2005dual also show limited adaptability to OOD scenarios.
J.1 AI and Algorithmic Decision-making in Emergency Response
In recent years, various AI-driven strategies have significantly advanced emergency response scenarios, particularly in the detection and localization of hazardous sources. Information-theoretic methods like Infotaxis vergassola2007infotaxis; masson2009chasing and Entrotaxis hutchinson2018entrotaxis optimize sensing by maximizing expected information gain or reducing entropy. Dual-control strategies such as DCEE explicitly balance exploitation (moving toward the best estimate) and exploration (reducing uncertainty), enabling effective autonomous searches in dynamic environments ferrari2020dual.
Complementary to planning-based methods, reinforcement learning (RL) techniques have also been successfully integrated into chemical source localization. Zhao et al. (2022) applied deep RL for plume tracing, directly mapping observations to navigation policies zhao2022deep. Park et al. (2022) developed GMM-PFRL, combining Gaussian mixture modeling with RL to enhance search efficiency park2022gaussian. Park and Kim (2020) employed adaptive Bayesian control techniques with RL strategies to dynamically guide sensors under uncertainty park2020adaptive.
J.2 Particle Filtering and Its Limitations in Emergency Response
Sequential Bayesian inference methods, especially particle filtering (PF), are widely adopted for source localization tasks under uncertainty ristic2004beyond; thrun2005probabilistic; arulampalam2002tutorial. Despite their flexibility, PFs operating under the stationary bootstrap baseline can exhibit the S-PSI effect: once the initial prior excludes the true state region, subsequent updates cannot recover it. Such rigidity significantly reduces practical effectiveness in disaster management where early information may be inaccurate thrun2005probabilistic. Classical perturbation remedies can alleviate S-PSI but, being always-on, risk over-diffusion and inefficiency.
J.3 Methods for Correcting Prior and Model Errors
To address these limitations, several approaches introduce adaptive exploration, e.g., injecting exploratory particles or using MCMC moves after resampling sarkka2013bayesian; liu2001combined. These methods expand support but often lack principled triggers or validation. Another line employs RL and adaptive control: AGDC uses PF-based uncertainty (variance or KL divergence) as intrinsic rewards shi2024autonomous, PC-DQN compresses PF belief into DQN features zhao2022deep, and GMM-PFRL fits mixture models to PF posterior park2022gaussian. Such methods highlight RL’s ability to adapt to prior mismatches, though they still rely on PF inference that may exhibit S-PSI under baseline assumptions.
J.4 Policy Robustness and Correction Strategies
Ensuring robustness of emergency response policies against early estimation errors is critical. Robust optimization and adaptive decision frameworks emphasize continuous revision of early assumptions kurniawati2008sarsop; pineau2006anytime. This parallels OOD detection in ML, where models identify distribution shifts and adjust silver2010monte. Such perspectives underscore the need for principled mechanisms to expand support when prior assumptions are violated.
J.5 Limitations of Baselines and Our Positioning
Despite progress, existing methods falter under severe prior misalignment. Information-driven planners (Infotaxis, Entrotaxis, DCEE) assume correct prior coverage; RL-based approaches (AGDC, PC-DQN, GMM-PFRL) degrade when true states lie outside initial support. Classical perturbations can relax S-PSI but lack control. Our work fills this gap: DEPF introduces belief-triggered exploratory injection, covariance-scaled diffusion, and MH validation, enabling minimal-bias, data-driven support expansion. This positions DEPF as a robust inference module for real-time emergency management under misaligned priors.
Appendix K Detailed Experimental Analysis
@@
K.1 Detailed analysis of Table 1 by scenario and scale
@@
Appendix L Hardware Usage
@@
Appendix M Resolving Stationarity–Induced Posterior–Support Lock–in: Theory and Clarifications
@@
Definition M.1 (S-PSI: Stationarity–Induced Posterior Support Invariance).
@@
Proposition M.2 (S-PSI under the stationary bootstrap baseline).
@@
Theorem M.3 (DEPF resolves S-PSI).
@@
Corollary M.4 (Finite-step support–recovery bound).
@@
M.2 Action Selection and the Role of the Controller
@@
M.3 Positioning w.r.t. Classical and Tempered/Bridge SMC
@@
M.4 Empirical Evidence in Support of the Theory
@@
Appendix N Theoretical Analysis: Stationarity-Induced Posterior Support Invariance (S-PSI) and Guarantees of DEPF
@@
N.1 S-PSI under the stationary bootstrap baseline
@@
Assumption N.1 (S0: zero transition, no rejuvenation).
@@
Definition N.2 (Stationarity-Induced Posterior Support Invariance (S-PSI)).
Under Assumption N.1, if particles are initialised in $S_{\text{prior}}$, then the posterior support remains confined to the initial prior support for all $k$.
Proposition N.3 (Baseline S-PSI).
Under Assumption N.1,
@@
S‑PSI is a baseline pathology of the stationary bootstrap with no rejuvenation, not an inherent limitation of particle filtering; classical perturbations (jittering/roughening/resample–move) can, in principle, relax it but operate in an always-on fashion and may be inefficient at scale (see main text §3.4/ §4 and §I). Gordon1993; Doucet2000
N.2 DEPF mechanisms (recap)
@@
N.4 Main guarantee: DEPF resolves S-PSI
Theorem N.4 (Asymptotic coverage under S-PSI).
Assume S‑PSI holds for the baseline (Def. N.2) and $\Theta^{\star}\notin S_{\text{prior}}$. Under (C1)(C2) and standard SMC regularity (bounded likelihood; stable tempering/entropy regularisation; KDE bandwidth $h_{\text{opt}}=A\,N^{-1/(n+4)}$), as $N\to\infty$ the DEPF support covers the true state with probability one:
@@
N.5 Entropy/tempering regularisation
@@
N.6 MH acceptance with symmetric Gaussian proposals
@@
N.7 Discussion: why DEPF breaks S-PSI reliably
Compared with always-on perturbations (jittering/roughening/resample–move), DEPF: (i) triggers support expansion only when belief–data inconsistency is detected, (ii) scales diffusion with the current covariance and a KDE bandwidth that vanishes at the right rate, and (iii) validates proposals by MH to preserve Bayesian coherence. These elements together provide both the asymptotic guarantee (Theorem N.4) and the finite-step bound (8).
Appendix O Belief-Aware RL Controller with Autonomous Goal Detection & Cessation
@@
O.1 POMDP Formalization and Belief State

### v3

@@ 1 Introduction
In disaster emergency management, early decisions play a pivotal role in shaping the outcome of a crisis Steel (2010); Kruschke (2010). An initial state estimate—for example, an early guess of a hazardous leak’s location—is often made with scant data under intense time pressure. Such early estimates carry great weight in guiding subsequent actions, yet they are highly vulnerable to error due to uncertainty and human subjective judgment. If the initial assumption is mistaken, the entire response effort can become locked into a false premise, failing to adjust even as new data arrives. This lock-in effect in high-stakes scenarios leads to catastrophic delays, misallocation of resources, and heightened risks to society. The fragility of such initial state assumptions thus poses a serious challenge for social decision-making processes, revealing the need for methods that can rapidly correct initial mistakes in real time.
@@
From a robotics and Bayesian inference perspective, the above scenario is essentially a state estimation problem: an agent must form an initial belief (prior) about the latent state (e.g., the hazard’s location) and update this belief as new sensor data arrives. If the initial belief is mis-specified, the goal is to correct this incorrect state estimate in real time by incorporating incoming evidence. For clarity, we use the term initial policy error to denote such a mistaken initial state estimate under prior uncertainty, which our approach aims to correct. In recent years, particle filtering (PF) has become a go-to approach for sequential Bayesian state estimation in these settings Ristic (2013); Gordon et al. (1993a); Doucet et al. (2001a). PF offers a principled way to update beliefs by fusing sensor data with prior knowledge, making it a representative framework to tackle the problem of erroneous initial state estimates in emergency response and other high-stakes decision domains. However, existing Bayesian filtering methods Fox et al. (2003); Smidl and Quinn (2008) struggle with erroneous initial state estimates, often failing to recover when the true state lies outside the initial belief. 1) Bootstrap particle filters Candy (2007); Gordon et al. (1995) tend to remain confined to the support of the initial prior, hindering exploration beyond the originally assumed region when operated under a stationary bootstrap baseline with zero transition and no rejuvenation. We formalize this baseline-specific lock-in as Stationarity-Induced Posterior Support Invariance (S-PSI): under the zero-transition, no-rejuvenation baseline, the posterior cannot escape the initial support. Importantly, S-PSI is not an inherent limitation of PF; classical countermeasures (e.g., jittering, roughening) can in principle expand support (we include these as baselines in our experiments). 2) More advanced PF variants,such as auxiliary particle filters Mountney et al. (2009); Branchini and Elvira (2021) or filters with optimal proposal distributions, can partially mitigate bias when the true state has a small non-zero prior probability Fox (2001); Douc and Cappé (2005); Liu and Chen (1998); Doucet et al. (2000); Arulampalam et al. (2002), but they fundamentally cannot handle cases where the true state was assigned zero initial probability. Under the standard Bayesian update, any region with zero prior mass will remain at zero posterior mass indefinitely, meaning the algorithm can never discover a completely excluded possibility under this baseline. 3) Attempts to address these issues by augmenting particle filters have had limited success. Some works inject noise or broaden the prior artificially, and others integrate reinforcement learning (RL) with PF Shi et al. (2024); Zhao et al. (2022); Park et al. (2022) to actively guide sensor exploration. While such approaches can improve data collection, they may inherit the same blind spots from a mis-specified prior and often introduce significant complexity and resource demands. Without a new perspective, Bayesian trackers and decision methods remain at risk of locking onto an incorrect initial belief, especially under the S-PSI.
@@ 2 Related Work
In emergency localization scenarios Wu et al. (2021); Hite (2019), Bayesian filtering Fox et al. (2003); Smidl and Quinn (2008); Quinlan and Middleton (2009) methods such as the bootstrap particle filter leverage Bayesian inference to iteratively update state estimates but typically assume correctly-specified initial priors, thus may become ineffective under severely misaligned early assumptions when operated under a stationary bootstrap baseline (zero transition, no rejuvenation), due to Stationarity-Induced Posterior Support Invariance (S-PSI) Gordon et al. (1995); Candy (2007). S-PSI is a baseline-specific diagnostic rather than an inherent limitation of PF; classical countermeasures such as jittering/roughening or resample–move can, in principle, expand support, and we include them as baselines in §5. Advanced particle filter variants, including auxiliary particle filters Mountney et al. (2009); Branchini and Elvira (2021) and filters using optimal proposal distributions, improve proposal quality and sample efficiency but still fail when the initial belief assigns zero prior probability to the true state Fox (2001); Arulampalam et al. (2002), i.e., in the presence of a zero-prior barrier without explicit support expansion. Meanwhile, information-theoretic methods like Infotaxis Vergassola et al. (2007), Entrotaxis Hutchinson et al. (2018) and DCEE Chen et al. (2021) focus sensor motions on maximizing expected information gain or reducing entropy, yet they typically operate within the belief support induced by the initial prior and thus cannot systematically correct severe prior misalignment in real time without a support-expanding inference layer. More recently, integrated reinforcement learning and particle filtering (RL-PF) methods have emerged—e.g., AGDC and its variants using KL-divergence or entropy-based intrinsic rewards Shi et al. (2024), PC-DQN Zhao et al. (2022), and GMM-PFRL Park et al. (2022). While these RL-driven methods exhibit stronger adaptive exploration, they can still inherit the zero-prior barrier when the underlying filtering layer does not expand support (cf. the S-PSI baseline). In contrast, our proposed DEPF explicitly addresses this gap by introducing a belief-triggered, validated support-expansion mechanism that operates at the inference layer, making it complementary to proposal-improvement filters, classical perturbations (jittering/roughening/rejuvenation), and planning/RL controllers, and yielding superior robustness under severely misaligned early assumptions.
@@ 3.5 POMDP formulation for belief‑aware sensor planning
The interaction between the mobile robot and the unknown gas plume can be framed as a partially observableMarkov decision process (POMDP) $\mathcal{M}=(\mathcal{S},\mathcal{A},\Omega,T,O,R,\gamma)$, where $\mathcal{S}$ is the latent state space containing the stationary source vector $\Theta=(x_{s},y_{s},q_{s},u_{s},\phi_{s},d_{s},\tau_{s})^{\top}$; $\mathcal{A}$ is the set of motion commands that move the robot in the plane; $\Omega$ is the observation space, where an observation at time $k$ is $o_{k}=(p_{k},z_{k})$ with $p_{k}\in\mathbb{R}^{2}$ the robot position and $z_{k}\in\mathbb{R}$ the sensor voltage. The transition kernel factorises as $T((p^{\prime},\Theta^{\prime})\mid(p,\Theta),a)=T_{p}(p^{\prime}\mid p,a)\,\delta(\Theta^{\prime}-\Theta)$: robot kinematics are deterministic while the source parameters remain unchanged. The observation model is $O(o_{k}\mid(p_{k},\Theta),a_{k-1})=p(z_{k}\mid p_{k},\Theta)$, where the likelihood $p(z_{k}\mid p_{k},\Theta)$ is the mixture‑Gaussian plume sensor model of §3.2. At decision time the agent cannot access $\Theta$ directly, so it reasons with the belief $b_{k}=p(\Theta\mid z_{1:k})$ supplied by the particle filter. We therefore feed the RL policy with the augmented information state $s^{\text{RL}}_{k}=(p_{k},z_{k},b_{k})$. The reward is chosen as the expected one‑step information gain $R_{k}=\mathbb{E}_{o_{k+1}}[D_{\mathrm{KL}}(b_{k+1}\,\|\,b_{k})]$, encouraging actions that shrink posterior uncertainty, and future rewards are discounted by $\gamma\in(0,1)$. The objective is to learn a policy $\pi^{\ast}$ that maximises the expected discounted return $J=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}R_{k+t}]$, thereby steering the robot along paths that are most informative about the hidden source. The details are provided in Appendix § N.
@@ 4 Diffusion-Driven Support Range Expansion
Diffusion-Driven Validation via MCMC: To ensure consistency with the target posterior distribution, a Metropolis-Hastings acceptance criterion Hastings (1970) validates the diffused particles. For each perturbed particle $\Theta_{k}^{(i)}$, the acceptance probability is: $\alpha_{i}=\frac{w_{\text{new}}^{(i)}}{w_{\text{old}}^{(i)}}\cdot\exp\left(-\frac{1}{2}\Delta\Theta_{k}^{(i)^{T}}\Sigma^{-1}\Delta\Theta_{k}^{(i)}\right).$ A uniformly sampled random variable $u_{i}\sim\mathcal{U}(0,1)$ determines whether the particle is accepted: $\Theta_{k}^{(i)}=\Theta_{k}^{(i)}-\Delta\Theta_{k}^{(i)}\cdot\mathbb{I}(\alpha_{i}<u_{i})$, where $\mathbb{I}(\alpha_{i}<u_{i})$ is the indicator function, which equals 1 when $\alpha_{i}<u_{i}$ is true, and 0 otherwise. This step ensures that the diffusion-driven expansion aligns with the posterior distribution, preserving the accuracy of the particle filter.
@@ 5.1 Evaluation Metrics and Baseline Algorithms
To evaluate our proposed approach and compare it against baseline algorithms, we use four distinct metrics: Operational Completion Efficacy (OCE), which measures how frequently emergency response missions meet their goals, with higher scores indicating better deployment effectiveness; Average Deployment Efficiency (ADE), representing the average distance traveled by response units, where shorter distances imply more efficient routing; Response Execution Velocity (REV), quantifying the time duration from deployment to task completion, with faster times signifying more efficient operations; and Localization Precision Score (LPS), assessing the accuracy of source localization by computing the average discrepancy between estimated and actual source locations, with smaller values denoting higher accuracy. Our proposed method is evaluated alongside various baseline algorithms, grouped according to their methodological foundations. The first group merges reinforcement learning with Bayesian inference and includes a single representative, AGDC Shi et al. (2024), which leverages the particle-filter RL posterior and uses intrinsic rewards derived from belief updates to guide exploration. Additionally, we include two other reinforcement learning approaches, PC-DQN Zhao et al. (2022) and GMM-PFRL Park et al. (2022), which independently leverage particle filtering parameters as states for RL training. The second group integrates planning and Bayesian inference approaches, represented by Infotaxis Vergassola et al. (2007), Entrotaxis Hutchinson et al. (2018), and DCEE Chen et al. (2021). Finally, we include a third group of classical SMC perturbation baselines that are not subject to S-PSI constraints: PF+Jittering Liu and Chen (1998); Doucet et al. (2000), PF+Roughening Gordon et al. (1993a; 1995), and PF+Rejuvenation Hastings (1970); Doucet et al. (2000). Implementation details and hyperparameter grids are given in §H and I.
@@ Appendix A Particle Filtering: Recursive Approximation of Posterior Distributions
Step 1. Sampling (Prediction): Gordon et al. (1993a) Particles are sampled from an importance distribution $q(\Theta_{k}\mid\Theta_{k-1},z_{k})$, typically approximated by the state transition model $P(\Theta_{k}\mid\Theta_{k-1})$. The propagation of particles is expressed as:
@@
Step 2. Weight Update: Doucet et al. (2001a) Particle weights are updated to reflect their likelihood given the new observation. The unnormalized weights are calculated as:
@@
Step 3. Normalization of Weights: Liu and Chen (1998) Weights are normalized to maintain their probabilistic interpretation:
@@
Step 4. Resampling: Doucet et al. (2000) To mitigate weight degeneracy (domination by a few particles), resampling occurs when the effective particle number $N_{\text{eff}}$ falls below a threshold:
@@ C.3 Entropy-Based Weight Smoothing
which can be viewed as Bayesian smoothing (Doucet et al., 2001b; Liu and Chen, 1998). In practice, we also adopt a standard tempering scheme
@@ C.4 Kernel-Induced Stochastic Diffusion
where $LL^{\top}=\Sigma$ is the Cholesky factorization of the weighted covariance. This is equivalent to a kernel density approximation (Silverman, 2018), and ensures that as $N\to\infty$ the perturbed empirical measure converges to the true posterior.
@@ C.5 Metropolis–Hastings Validation
This guarantees detailed balance and convergence to the true posterior distribution (Hastings, 1970).
@@ I.3 Pseudocode
Appendix J Detailed Experimental Analysis
@@
J.1 Detailed analysis of Table 1 by scenario and scale
@@
Appendix K Hardware Usage
@@
Appendix L Resolving Stationarity–Induced Posterior–Support Lock–in: Theory and Clarifications
@@
Definition L.1 (S-PSI: Stationarity–Induced Posterior Support Invariance).
@@
Proposition L.2 (S-PSI under the stationary bootstrap baseline).
@@
Theorem L.3 (DEPF resolves S-PSI).
@@
Corollary L.4 (Finite-step support–recovery bound).
@@
L.2 Action Selection and the Role of the Controller
@@
L.3 Positioning w.r.t. Classical and Tempered/Bridge SMC
@@
L.4 Empirical Evidence in Support of the Theory
@@
Appendix M Theoretical Analysis: Stationarity-Induced Posterior Support Invariance (S-PSI) and Guarantees of DEPF
@@
M.1 S-PSI under the stationary bootstrap baseline
@@
Assumption M.1 (S0: zero transition, no rejuvenation).
@@
Definition M.2 (Stationarity-Induced Posterior Support Invariance (S-PSI)).
Under Assumption M.1, if particles are initialised in $S_{\text{prior}}$, then the posterior support remains confined to the initial prior support for all $k$.
Proposition M.3 (Baseline S-PSI).
Under Assumption M.1,
@@
S‑PSI is a baseline pathology of the stationary bootstrap with no rejuvenation, not an inherent limitation of particle filtering; classical perturbations (jittering/roughening/resample–move) can, in principle, relax it but operate in an always-on fashion and may be inefficient at scale (see main text §3.4/ §4 and §I). Gordon et al. (1993b); Doucet2000
M.2 DEPF mechanisms (recap)
@@
M.4 Main guarantee: DEPF resolves S-PSI
Theorem M.4 (Asymptotic coverage under S-PSI).
Assume S‑PSI holds for the baseline (Def. M.2) and $\Theta^{\star}\notin S_{\text{prior}}$. Under (C1)(C2) and standard SMC regularity (bounded likelihood; stable tempering/entropy regularisation; KDE bandwidth $h_{\text{opt}}=A\,N^{-1/(n+4)}$), as $N\to\infty$ the DEPF support covers the true state with probability one:
@@
M.5 Entropy/tempering regularisation
@@
M.6 MH acceptance with symmetric Gaussian proposals
@@
M.7 Discussion: why DEPF breaks S-PSI reliably
Compared with always-on perturbations (jittering/roughening/resample–move), DEPF: (i) triggers support expansion only when belief–data inconsistency is detected, (ii) scales diffusion with the current covariance and a KDE bandwidth that vanishes at the right rate, and (iii) validates proposals by MH to preserve Bayesian coherence. These elements together provide both the asymptotic guarantee (Theorem M.4) and the finite-step bound (8).
Appendix N Belief-Aware RL Controller with Autonomous Goal Detection & Cessation
@@
N.1 POMDP Formalization and Belief State